OpenStudy (anonymous):

Please HELP! Factor this expression completely. x^3 + 1/8

OpenStudy (anonymous):

Use the formula \[ a^3+b^3=(a+b) \left(a^2-a b+b^2\right) \]

OpenStudy (anonymous):

a=x b=1/2

OpenStudy (anonymous):

Im not sure

OpenStudy (mathmale):

Hint: Look up "special products and factoring" in your textbook or online. Note that x^3 - 1/8 can be written as \[x^3+(1/2)^3.\] which has the same form as the (a^3+b^3) mentioned by eliassaab.

OpenStudy (anonymous):

so would it be (x + 1/2) ( x^2- x1/2 + 1/2^2) ??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[ \left(x+\frac{1}{2}\right) \left(x^2-\frac{x}{2}+\frac {1}{4}\right) \]

OpenStudy (mathmale):

Nice work, Power2Knowledge!!

OpenStudy (anonymous):

thanks